# D4C Glove-train: Solving the RPM and Bongard-logo Problem by Circumscribing and Building Distribution for Concepts

### v5

@@ VIII-D Ablation experiment on D2C
ResNet50 92.0 88.9 98.1 76.1 76.0
Triple-CFN 93.2 92.0 99.2 80.8 79.1
ResNet50+D2C 92.4 90.7 99.0 77.0 77.5
Triple-CFN+D2C 96.5 94.8 99.4 79.5 81.6
ResNet50+D2C- 93.5 84.2 99.5 77.0 76.2
Triple-CFN+D2C- 95.7 93.2 98.9 79.3 80.9
It can be observed from the experiment results recorded in Table VI that introducing generated samples as negative examples is difficult to fundamentally alter the network’s performance, as such behavior struggles to reorganize the concepts. In other words, D2C is restructuring the conceptual space, and the new concept space is highly coupled with the image style, with mutual exclusion between different cases of concepts. When the images generated by DDPM are introduced as auxiliary images into the Bongard-Logo case, they cannot achieve the above goal. Because D2C does not stop at the surface, that is, adding more auxiliary images to the Bongard-Logo case.
@@ IX Conclusion
D4C builds upon the principles of both D2C and D3C, exploring the roles of positive and negative examples in abstract reasoning problems from a perspective on distribution. Specifically, positive examples are utilized to depict the mean for deep networks, while negative examples contribute to describing the variance. Based on this understanding, D4C employs an adversarial approach, pitting the generator against the discriminator in repeated challenges. The generator continually produces negative examples that push the boundaries of the problem, while the discriminator refines its skills through repeated defense.
This iterative process not only enhances the discriminator’s ability to distinguish between genuine and generated examples but also improves the generator’s capacity to generate more sophisticated negative examples. Ultimately, this adversarial framework promotes the evolution of both components, leading to more robust and reliable solutions in abstract reasoning tasks.

### v9

@@ VIII-D Ablation experiment on D2C and D4C
ResNet50 (infoNCE baseline) 92.0 88.9 98.0 76.1 76.0
Triple-CFN 93.2 92.0 98.2 78.0 78.1
ResNet50+D2C- 91.6 90.0 97.9 76.3 76.1
Triple-CFN+D2C- 93.0 92.1 98.2 77.9 78.2
It can be observed from the experiment results recorded in Table VII that introducing generated samples as negative examples is difficult to fundamentally alter the network’s performance, as such behavior struggles to re-delineate the concepts. DDPM-generated data, sharing the same distribution as real data but with inherent uncertainties of concept and attribution, are limited in contributing to the delineation of concept boundaries.
D4C posits that the accuracy achieved by the discriminator in abstract reasoning problems is closely related to the quality of the positive and negative examples set in the task. The established examples serve to delineate the concept boundaries for abstract reasoning challenges. D2C, while attempting to define new concept boundaries for such problems, represents an imperfect methodology. Conversely, D4C emerges as a relatively more comprehensive approach. This ablation experiment aims to verify that, within the context of Bongard-Logo problems, the provision of meaningless negative examples—examples that do not meaningfully contribute to defining the boundaries of reasoning tasks—fails to enhance the discriminator’s effectiveness. This finding echoes similar observations made in the RAVEN dataset, reinforcing the significance and rationale behind the design of D4C. Furthermore, it reaffirms D4C’s assertion regarding the importance of high-quality positive and negative examples in abstract reasoning tasks. It also provides indirect evidence that the approaches of D2C and D4C extend beyond merely introducing additional negative examples. Instead, they aim to introduce negative examples that better delineate the boundaries of the problem.
@@ IX Conclusion
D4C builds upon the principles of both D2C and D3C, exploring the roles of positive and negative examples in abstract reasoning problems from a perspective on distribution. Specifically, positive examples are utilized to depict the mean for deep networks, while negative examples contribute to describing the variance. Based on this understanding, D4C employs an adversarial approach, pitting the generator against the discriminator in repeated challenges. The generator continually produces negative examples that push the boundaries of the problem, while the discriminator refines its skills through repeated defense. This iterative process not only enhances the discriminator’s ability to distinguish between positive and negative examples but also improves the generator’s capacity to generate more sophisticated negative examples. Ultimately, this adversarial framework promotes the evolution of both components, leading to more robust and reliable solutions in abstract reasoning tasks.
For future, the D4C method presents a novel approach for abstract reasoning problems and even other discriminative tasks. Its exploration of the identities of positive and negative examples may provide insights for the development of other works.